The age of Tim is 1 year more than twice of Carla's age. Three years ago, he was thrice as old as she was then. Find their present ages.


The sum of Gelard's age and Fred's is 53 an one-half ( 1/2 ) as old as Fred. Then How old is Fred?

Answer by thinktank.experts(5)

let carla age =x yrs.
tim=2x+1
acc to the problem---
2x+1-3=3(x-3)
2x-2=3x-9
-x=-7
so, x=7
carla age=7yrs
tim age=2*7+1=15yrs..